MsBrains (ms-brains):

Solve the equation. 6=a/4+2

MsBrains (ms-brains):

Can someone show me how to do this, so I can learn?

OpenStudy (anonymous):

I can help. @Ms-Brains

MsBrains (ms-brains):

Thanks

OpenStudy (anonymous):

Since a is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. \[\frac{ a }{ 4 }+2=6\] Then move all the terms that don't contain a to the right had side of this equation. \[\frac{ a }{ 4 }=-2+6\] Add -2 and 6 to get 4. \[\frac{ a }{ 4 }=4\] Multiply both sides of the the equation by 4. \[a= 4 \times (4)\] Then Simplify the right hand side. \[a =16\] Then plug in the 16 to your equation for a. \[6= \frac{ 16 }{ 4 }+2\] 16 divided by 4 equals 4. \[6= 4 + 2\] Then 6=6 @Ms-Brains

OpenStudy (anonymous):

Did you understand this? @Ms-Brains

OpenStudy (anonymous):

Or you could of thought of this question as simple math. 4x4= 16 16 divide 4= 4

MsBrains (ms-brains):

Yes, I get it! Thank you Austin!

OpenStudy (anonymous):

I can help with more questions if you need the help. @Ms-Brains

OpenStudy (mathmale):

Nice work, Austin. But please don't do all the work for others; get them involved in finding their own solutions, please.